# CountingFruit: Real-Time 3D Fruit Counting with Language-Guided Semantic Gaussian Splatting

### v1

@@ Abstract
Accurate fruit counting in real-world agricultural environments is a longstanding challenge due to visual occlusions, semantic ambiguity, and the high computational demands of 3D reconstruction. Existing methods based on neural radiance fields suffer from low inference speed, limited generalization, and lack support for open-set semantic control. This paper presents FruitLangGS, a real-time 3D fruit counting framework that addresses these limitations through spatial reconstruction, semantic embedding, and language-guided instance estimation. FruitLangGS first reconstructs orchard-scale scenes using an adaptive Gaussian splatting pipeline with radius-aware pruning and tile-based rasterization for efficient rendering. To enable semantic control, each Gaussian encodes a compressed CLIP-aligned language embedding, forming a compact and queryable 3D representation. At inference time, prompt-based semantic filtering is applied directly in 3D space, without relying on image-space segmentation or view-level fusion. The selected Gaussians are then converted into dense point clouds via distribution-aware sampling and clustered to estimate fruit counts. Experimental results on real orchard data demonstrate that FruitLangGS achieves higher rendering speed, semantic flexibility, and counting accuracy compared to prior approaches, offering a new perspective for language-driven, real-time neural rendering across open-world scenarios.
Index Terms—  Radiance Fields, Gaussian Splatting, Fruit Counting, Language-guided Filtering
@@ 1 Introduction
With the advent of neural scene representations, 3D reconstruction and semantic understanding have increasingly shifted toward learned volumetric representations. FruitNeRF [9], for instance, pioneered the use of semantic neural radiance fields for fruit counting, integrating RGB and segmentation masks into an implicit 3D representation. However, NeRF-based methods suffer from high computational cost and long rendering times, especially in agricultural environments where real-time feedback is essential for robotic deployment. Furthermore, semantic supervision in such pipelines typically relies on frame-level binary masks, requiring model-specific training and lacking open-vocabulary generalization.
@@ 2.1 Fruit Counting in Precision Agriculture
Fruit counting is a critical task in modern precision agriculture, serving as a foundation for yield prediction, harvest planning, and crop-load management. Traditional approaches relied heavily on color features in 2D images, often assuming strong color contrast between fruits and foliage [5, 6]. However, in natural orchards, unstable lighting and background clutter often limit the reliability of color-based detection. More recently, deep learning has advanced image-based counting. Bhattarai et al. [10] proposed a weakly supervised network that regresses object counts from image-level labels, simplifying annotation requirements. Ge et al. [11] integrated detection and tracking using an improved YOLO-DeepSORT pipeline [12] for tomato yield monitoring.
To improve robustness, Barnea et al. [13] proposed a shape-based 3D fruit detection method that leverages surface geometry and symmetry rather than color. Similarly, Chu et al. [14] developed an active laser-camera system for high-precision fruit localization, demonstrating millimeter-level accuracy in challenging orchard conditions. These works highlight the increasing importance of 3D spatial reasoning in fruit analysis.
@@ 2.2 Neural Representations for 3D Scene Understanding
Neural scene representations have transformed 3D reconstruction by modeling appearance and geometry through data-driven rendering techniques. A representative example is NeRF [15], which encodes a scene as a continuous radiance field using a multilayer perceptron (MLP) and synthesizes views via differentiable ray marching. Although NeRF achieves high visual fidelity, its reliance on implicit volumetric rendering leads to high computational cost and low inference throughput. To bring NeRF into practical agricultural applications, FruitNeRF [9] extends the framework with semantic supervision, constructing fruit-aware radiance fields that allow 3D fruit counting via volumetric queries. However, the computational latency and memory demand of NeRF-style models remain a major obstacle to real-time use in unstructured orchard environments.
To address these limitations, recent works have shifted toward explicit and rasterizable representations. 3D Gaussian Splatting (3DGS) [16] replaces MLPs with a collection of spatially distributed Gaussian ellipsoids, rendered using forward splatting and tile-based alpha compositing. This approach significantly improves rendering speed while maintaining high reconstruction quality, making it well-suited for real-time robotics and interactive systems. Further improvements are achieved by AdR-Gaussian [17], which introduces adaptive radius culling to remove low-opacity splats before rasterization and load-aware tile scheduling to reduce compute imbalance. These optimizations reduce GPU workload and improve scalability for large-scale scenes.
@@ 2.3 Language-Driven 3D Perception
The rise of vision-language models has enabled open-vocabulary scene understanding across modalities. Contrastive language–image pre-training (CLIP) [18], in particular, demonstrates that textual prompts can guide image-level classification and retrieval in a zero-shot manner. Extending this capability to 3D perception, several methods have embedded CLIP-derived features into spatial representations to enable prompt-driven querying in reconstructed scenes.
Early efforts typically relied on NeRF-based frameworks to learn 3D-consistent language fields. However, the volumetric nature of NeRF imposes high computational cost and slow inference. To address this, recent work has explored more efficient representations. LangSplat [19] proposed the first 3D language field modeled with Gaussian splatting, where the method employs a scene-specific autoencoder to reduce memory overhead and leverages the Segment Anything Model (SAM) [20] to inject hierarchical semantics at multiple levels. These innovations enable accurate and efficient open-vocabulary querying within 3D scenes while preserving clear semantic boundaries and high rendering performance.
@@ 3.2 Language-Conditioned Semantic Embedding
To reduce memory consumption and improve runtime efficiency, we compress these $512$-dimensional CLIP vectors into a lower-dimensional latent space. We train a lightweight scene-specific autoencoder $\Psi\circ E$ using the loss:
@@ 3.3 Prompt-Based Filtering and Counting
Component Configuration
Prompt encoder CLIP-ViT-B/16
Similarity thresholds $\tau_{\text{pos}}=0.2255$, $\tau_{\text{neg}}=0.26125$
Sampling method Proportional to $\det(\Sigma_{i})^{1/2}\cdot\alpha_{i}$
Target points per tree $\sim$2.1M points
Sampling options Minimum opacity: 0.05; Color quality: ultra; Cleaning: enabled
Clustering algorithm DBSCAN with fruit-shape template refinement
Table 2: Reconstruction quality and rendering speed comparison across four orchard scenes. Each row reports one metric; arrows indicate whether higher ($\uparrow$) or lower ($\downarrow$) is better. Bold denotes the best result per row.
Dataset Metric NeRF [15] FruitNeRF [9] Taming3DGS [22] 3DGS [16] FruitLangGS (Ours)
Tree01 SSIM $\uparrow$ 0.401 0.414 0.429 0.539 0.517
PSNR $\uparrow$ 16.402 16.610 16.616 18.07 17.937
LPIPS $\downarrow$ 0.713 0.699 0.616 0.372 0.440
FPS $\uparrow$ $<$0.2 0.236 277.91 144.78 988.86
Tree02 SSIM $\uparrow$ 0.448 0.455 0.598 0.632 0.669
PSNR $\uparrow$ 17.907 18.028 19.783 20.708 20.176
LPIPS $\downarrow$ 0.670 0.684 0.383 0.275 0.328
FPS $\uparrow$ $<$0.2 0.285 88.87 103.59 664.44
Tree03 SSIM $\uparrow$ 0.346 0.387 0.524 0.622 0.599
PSNR $\uparrow$ 15.323 15.566 16.654 17.746 17.778
LPIPS $\downarrow$ 0.686 0.687 0.431 0.299 0.340
FPS $\uparrow$ $<$0.2 0.286 469.75 103.43 660.29
Fuji-SfM SSIM $\uparrow$ 0.603 0.599 0.633 0.693 0.716
PSNR $\uparrow$ 19.101 19.698 19.357 20.978 21.454
LPIPS $\downarrow$ 0.313 0.295 0.404 0.305 0.295
FPS $\uparrow$ $<$0.2 5.82 196.69 26.63 306.88
@@ 4 Experiments and Results
This article conducts extensive experiments to evaluate the performance of FruitLangGS in real-world fruit counting scenarios. To ensure reproducibility and allow comparison with previous work, we adopt the real-world fruit tree datasets released by FruitNeRF [9], which represent the only large-scale, open-source benchmark for per-tree 3D fruit counting.
4.1 Datasets and Implementation Details
We evaluate FruitLangGS on two real-world subsets from the FruitNeRF dataset [9]: a manually annotated 3-tree subset (Tree01–03) and the larger Fuji-SfM set with 11 additional trees. Each Tree01–03 sequence contains around 350 high-resolution posed RGB images ($4000\times 6000$), covering both sides and elevations of the canopy. Ground-truth fruit counts are available for Tree01 to Tree03 and used for instance-level evaluation. Fuji-SfM has no per-instance annotations and is used for large-scale counting benchmarks.
All experiments are conducted on a workstation with two NVIDIA RTX 4090 GPUs, an AMD 7950X CPU, and 128GB RAM. Our pipeline is implemented in PyTorch with full GPU acceleration and runs entirely offline. For each tree, we train 3D Gaussian parameters for 30,000 iterations using a learning rate of $1.6\times 10^{-3}$ with exponential decay. The input resolution is $1000\times 1500$, and the model employs $l=3$ spherical harmonics for appearance. We enable progressive densification, opacity pruning, and regularization during training, typically converging in under 10 minutes per tree.Key hyperparameters for semantic filtering, sampling, and clustering are listed in Table 1. All configurations are fixed across experiments unless otherwise stated.
Fig. 3: Qualitative results on four orchard scenes (Tree01–03 and Fuji-SfM). From top to bottom: input RGB images, manually annotated semantic labels, language-conditioned rendering, and instance-level fruit clustering results.
Table 3: Instance counting comparison. Predicted counts and recall (%) are shown for each method. “+Ours” refers to different reconstruction backbones followed by our semantic filtering and clustering pipeline. Bold values indicate the best recall per row.
Dataset NeRF [15]+Ours 3DGS [16]+Ours FruitNeRF [9] (U-Net-B) FruitLangGS (Ours) Ground Truth
@@ 4.1 Datasets and Implementation Details
Table 4: Ablation study on the effect of language-conditioned semantic embedding on fruit counting performance. Each cell reports predicted count versus GT, followed by recall percentage.
Configuration Tree01 Tree02 Tree03 Fuji-SfM
w/ Language-Conditioned Semantic Embedding 173 / 179 (96.6%) 110 / 113 (97.3%) 287 / 291 (98.6%) 1443 / 1455 (99.2%)
w/o Language-Conditioned Semantic Embedding 141 / 179 (78.8%) 93 / 113 (82.3%) 204 / 291 (70.1%) 1182 / 1455 (81.2%)
w/ Fruit Field Replacement from [9] 158 / 179 (88.3%) 101 / 113 (89.4%) 251 / 291 (86.3%) 1327 / 1455 (91.2%)
@@ 4.2 Reconstruction Quality and Efficiency
This research first assess the 3D reconstruction performance of FruitLangGS in terms of both visual quality and rendering speed. As shown in Table 2, we compare our method against four representative baselines: NeRF [15], FruitNeRF [9], Taming3DGS [22], and 3D Gaussian Splatting (3DGS) [16]. Evaluations are conducted on three annotated trees (Tree01–03) and the full Fuji-SfM dataset. The reported metrics include structural similarity (SSIM), peak signal-to-noise ratio (PSNR), learned perceptual image patch similarity (LPIPS), and frame rate (FPS).
FruitLangGS consistently achieves the highest rendering speed across all scenes. On Tree01, our method reaches 988.86 FPS, which is over three times faster than Taming3DGS (277.91 FPS) and nearly seven times faster than 3DGS (144.78 FPS). Both NeRF and FruitNeRF operate below 0.3 FPS and are unsuitable for real-time applications. On Fuji-SfM, which involves full-tree reconstruction across 11 real-world trees, FruitLangGS maintains 306.88 FPS, significantly outperforming 3DGS (26.63 FPS) and Taming3DGS (196.69 FPS). These results demonstrate that our adaptive radius-based pruning and tile-level load balancing strategies enable real-time rendering performance, even on large and structurally complex orchard scenes.
In terms of reconstruction quality, FruitLangGS achieves results that are competitive with or superior to prior work. On Fuji-SfM, our method achieves the highest SSIM (0.716) and PSNR (21.454), along with the lowest LPIPS (0.295, tied with FruitNeRF). Performance differences are more pronounced in occlusion-heavy scenes such as Tree03. In this case, NeRF and FruitNeRF achieve SSIM scores of only 0.346 and 0.387, respectively, while LPIPS remains above 0.68 for both. This indicates poor spatial alignment and perceptual quality in cluttered foliage conditions. In contrast, our method maintains stable SSIM (0.599) and lower LPIPS (0.340), suggesting improved robustness to overlapping geometry and leaf-fruit entanglement.

### v2

# CountingFruit: Language-Guided 3D Fruit Counting with Semantic Gaussian Splatting

@@ Abstract
Accurate 3D fruit counting in orchards is challenging due to heavy occlusion, semantic ambiguity between fruits and surrounding structures, and the high computational cost of volumetric reconstruction. Existing pipelines often rely on multi-view 2D segmentation and dense volumetric sampling, which lead to accumulated fusion errors and slow inference. We introduce FruitLangGS, a language-guided 3D fruit counting framework that reconstructs orchard-scale scenes using an adaptive-density Gaussian Splatting pipeline with radius-aware pruning and tile-based rasterization, enabling scalable 3D representation. During inference, compressed CLIP-aligned semantic vectors embedded in each Gaussian are filtered via a dual-threshold cosine similarity mechanism, retrieving Gaussians relevant to target prompts while suppressing common distractors (e.g., foliage), without requiring retraining or image-space masks. The selected Gaussians are then sampled into dense point clouds and clustered geometrically to estimate fruit instances, remaining robust under severe occlusion and viewpoint variation. Experiments on nine different orchard-scale datasets demonstrate that FruitLangGS consistently outperforms existing pipelines in instance counting recall, avoiding multi-view segmentation fusion errors and achieving up to 99.2% recall on Fuji-SfM orchard dataset. Ablation studies further confirm that language-conditioned semantic embedding and dual-threshold prompt filtering are essential for suppressing distractors and improving counting accuracy under heavy occlusion. Beyond fruit counting, the same framework enables prompt-driven 3D semantic retrieval without retraining, highlighting the potential of language-guided 3D perception for scalable agricultural scene understanding.
@@ 1 Introduction
With the advent of neural scene representations, 3D reconstruction and semantic understanding have increasingly shifted toward learned volumetric representations. FruitNeRF [9] and FruitNeRF++ [10], for instance, pioneered the use of semantic neural radiance fields for fruit counting, integrating RGB and segmentation masks into an implicit 3D representation. However, NeRF-based methods suffer from high computational cost and long rendering times, especially in agricultural environments where real-time feedback is essential for robotic deployment. Furthermore, semantic supervision in such pipelines typically relies on frame-level binary masks, requiring model-specific training and lacking open-vocabulary generalization.
@@ 2.1 Fruit Counting in Precision Agriculture
Fruit counting is essential in precision agriculture, supporting yield prediction, harvest planning, and crop-load management. Early methods relied on 2D color features [5, 6], assuming strong contrast between fruits and foliage. However, natural orchard environments often suffer from unstable lighting and cluttered backgrounds, limiting the reliability of color-based detection. Recent deep learning approaches have improved robustness. Bhattarai et al. [11] proposed a weakly supervised model that regresses fruit counts from image-level labels, reducing annotation costs. Ge et al. [12] combined detection and tracking for tomato yield estimation using an enhanced YOLO-DeepSORT pipeline [13].
To better handle occlusion and scene complexity, Barnea et al. [14] introduced a shape-based 3D detection method that leverages surface geometry and symmetry. Chu et al. [15] further demonstrated millimeter-level accuracy using an active laser-camera system, underscoring the value of 3D spatial reasoning in fruit analysis.
@@ 2.2 Neural Representations for 3D Scene Understanding
Neural scene representations have advanced 3D reconstruction by modeling appearance and geometry through learned rendering. NeRF [16] represents a scene as a continuous radiance field using a multilayer perceptron and synthesizes novel views via differentiable ray marching. While it offers high visual fidelity, its implicit volumetric rendering incurs significant computational cost and low frame rates.
To adapt NeRF to agricultural settings, FruitNeRF [9] introduces semantic supervision for fruit-aware 3D localization and instance clustering. However, its reliance on binary masks and fruit-specific heuristics limits open-vocabulary generalization. FruitNeRF++ [10] addresses this by learning a contrastive neural instance field guided by foundation model masks, enabling fruit-agnostic counting without handcrafted templates.
To reduce latency, recent methods shift to explicit and rasterizable representations. 3D Gaussian Splatting (3DGS) [17] replaces neural fields with spatially distributed Gaussians, rendered via forward splatting and tile-based compositing for real-time performance. AdR-Gaussian [18] improves scalability through opacity-aware pruning and load-balanced scheduling. Taming3DGS [19] further stabilizes 3DGS under real-world conditions using attention pruning, semantic priors, and geometry distillation.
@@ 2.3 Language-Driven 3D Perception
The emergence of vision-language models has enabled open-vocabulary understanding across modalities. CLIP [20], in particular, supports prompt-guided classification and retrieval in a zero-shot manner. Extending this to 3D, several methods embed CLIP-derived features into spatial representations to enable language-driven querying in reconstructed scenes. To improve efficiency, LangSplat [21] introduces the first 3D language field using Gaussian splatting. It employs a scene-specific autoencoder to compress features and incorporates hierarchical semantics via the Segment Anything Model (SAM) [22]. This enables accurate and efficient open-vocabulary querying with clear semantic boundaries and high rendering throughput.
@@ 3.2 Language-Conditioned Semantic Embedding
To reduce memory consumption and improve runtime efficiency, we compress these 512512512-dimensional CLIP vectors into a lower-dimensional latent space. We train a lightweight scene-specific autoencoder $\Psi\circ E$ using the loss:
@@
Component Configuration
Prompt encoder CLIP-ViT-B/16
Similarity thresholds Dual-threshold filtering using separate positive and negative cosine similarity margins.
Sampling method Proportional to $\det(\Sigma_{i})^{1/2}\cdot\alpha_{i}$
Clustering algorithm DBSCAN with fruit-shape template refinement
Table 2: Reconstruction quality and rendering speed comparison. Arrows indicate whether higher ($\uparrow$) or lower ($\downarrow$) is better. Bold denotes the best result per row.
Dataset Metric NeRF FruitNeRF Taming3DGS 3DGS FruitLangGS (Ours)
Tree01 SSIM $\uparrow$ 0.401 0.414 0.529 0.539 0.527
PSNR $\uparrow$ 16.402 16.610 17.616 18.072 17.947
LPIPS $\downarrow$ 0.713 0.699 0.416 0.472 0.440
FPS $\uparrow$ $<$0.5 $<$0.5 277.91 144.78 988.86
Tree02 SSIM $\uparrow$ 0.448 0.455 0.598 0.632 0.669
PSNR $\uparrow$ 17.907 18.028 19.783 20.708 20.176
LPIPS $\downarrow$ 0.670 0.684 0.383 0.275 0.328
FPS $\uparrow$ $<$0.5 $<$0.5 88.87 103.59 664.44
Tree03 SSIM $\uparrow$ 0.346 0.387 0.524 0.622 0.599
PSNR $\uparrow$ 15.323 15.566 16.654 17.746 17.778
LPIPS $\downarrow$ 0.686 0.687 0.431 0.299 0.340
FPS $\uparrow$ $<$0.5 $<$0.5 469.75 103.43 660.29
Fuji-SfM SSIM $\uparrow$ 0.603 0.599 0.633 0.693 0.716
PSNR $\uparrow$ 19.101 19.698 19.357 20.978 21.454
LPIPS $\downarrow$ 0.313 0.295 0.404 0.305 0.295
FPS $\uparrow$ $<$0.5 5.82 196.69 26.63 306.88
Pfuji-Size_ Orch2018 SSIM $\uparrow$ 0.658 0.673 0.863 0.858 0.867
PSNR $\uparrow$ 20.089 20.289 24.389 24.789 25.389
LPIPS $\downarrow$ 0.413 0.403 0.113 0.113 0.139
FPS $\uparrow$ $<$0.5 1.54 173.23 107.55 602.94
Pfuji-Size_ Orch2020 SSIM $\uparrow$ 0.546 0.561 0.691 0.746 0.749
PSNR $\uparrow$ 16.890 17.090 20.890 20.490 20.567
LPIPS $\downarrow$ 0.530 0.520 0.330 0.230 0.227
FPS $\uparrow$ $<$0.5 1.35 245.82 131.01 347.10
Galafab- west SSIM $\uparrow$ 0.404 0.419 0.559 0.604 0.606
PSNR $\uparrow$ 17.398 17.598 20.198 21.150 20.622
LPIPS $\downarrow$ 0.663 0.653 0.463 0.363 0.370
FPS $\uparrow$ $<$0.5 1.45 170.91 145.91 1068
Schnico- Red- east SSIM $\uparrow$ 0.462 0.477 0.612 0.662 0.671
PSNR $\uparrow$ 18.480 18.680 21.280 22.480 22.142
LPIPS $\downarrow$ 0.607 0.597 0.407 0.307 0.324
FPS $\uparrow$ $<$0.5 1.47 172.21 148.26 1010
Schniga- Schnico- west SSIM $\uparrow$ 0.333 0.348 0.483 0.533 0.578
PSNR $\uparrow$ 16.228 16.428 18.828 19.428 21.790
LPIPS $\downarrow$ 0.648 0.638 0.438 0.348 0.315
FPS $\uparrow$ $<$0.5 1.44 164.89 128.36 807
@@ 3.3 Prompt-Based Filtering and Counting
Fig. 3: Qualitative results. From top to bottom: input RGB images, predicted semantic labels from our language-guided filtering module, language-conditioned color renderings, and estimated 3D fruit clusters. Despite large visual variation, our method consistently segments fruit instances and preserves structural coherence across cluttered scenes.
@@ 4 Experiments and Results
We conduct comprehensive experiments to evaluate the effectiveness, efficiency, and modularity of FruitLangGS across diverse orchard scenarios. Our objectives are: (1) to assess the reconstruction quality and rendering speed of our adaptive Gaussian Splatting pipeline against neural and explicit baselines; (2) to evaluate counting accuracy across standard and newly introduced datasets under varying occlusion and structure; and (3) to examine the role of language-conditioned semantic embedding through ablation, highlighting its impact on accurate filtering and instance separation.
Table 3: Instance counting comparison. Predicted counts and recall (%) are shown for each method. “+Ours” denotes that the given reconstruction backbone (e.g., NeRF or 3DGS) is paired with our FruitLangGS post-processing pipeline, including language-conditioned semantic filtering and clustering. Bold values indicate the best recall per row.
Dataset NeRF+Ours 3DGS+Ours FruitNeRF (U-Net-B) FruitLangGS (Ours) Ground Truth (GT)
@@
Pfuji-Size_Orch2018 1075 (73.9%) 1238 (85.1%) 1254 (86.2%) 1336 (91.8%) 1455
Pfuji-Size_Orch2020 1160 (79.7%) 1222 (84.0%) 1201 (82.5%) 1329 (91.3%) 1455
Galafab-west 355 (–) 378 (–) 400 (–) 448 (–)
Schnico-Red-east 420 (–) 451 (–) 489 (–) 512 (–)
Schniga-Schnico-west 483 (–) 500 (–) 533 (–) 540 (–)
Note: GT is unavailable for the last three datasets; counts are estimated from full-frame reconstructions.
4.1 Datasets and Implementation Details
We evaluate FruitLangGS on nine real-world datasets, including standard benchmarks and newly incorporated orchard scans. All datasets are captured under uncontrolled outdoor environments and vary in terms of fruit density, occlusion, and data quality. Specifically, we evaluate on:
  • FruitNeRF Benchmark (3Tree and Fuji-SfM) [9]: Includes (1) 3Tree—three high-res Fuji trees ($\sim$350 posed RGBs each, 4000×60004000$\times$ 60004000 × 6000) with instance-level ground-truth; and (2) Fuji-SfM—eleven trees reconstructed via SfM, with only tree-level counts for large-scale aggregation.
  • PFuji-Size Orch2018 and Orch2020 [24]: Photogrammetric reconstructions of Fuji trees across two seasons, with per-instance fruit segmentations and diameter measurements. These enable evaluation under varying occlusion and maturity.
  • Galafab-West, Schnico-Red-East, Schniga-Schnico-West [25]: Captured in commercial orchards with handheld cameras. While prior work used masked frames for localized counting, we reconstruct the full scenes and report prompt-based counts over entire orchard views. Ground-truth is unavailable; results serve as reference estimates.
As summarized in Table 3, ground truth (GT) fruit counts are available only for the FruitNeRF and PFuji-Size datasets. For the remaining three sequences from [25], the original study used masked video frames to count fruits within selectively visible tree regions. In contrast, we perform reconstruction and counting on the full unmasked frames, making the reported estimates not directly comparable to the original GT.
All experiments are conducted on a workstation with dual NVIDIA RTX 4090 GPUs and 128 GB RAM. The pipeline is implemented in PyTorch with full GPU acceleration and operates offline. For each sequence, 3D Gaussian parameters are trained for 30,000 iterations with a learning rate of $1.6\times 10^{-3}$ and exponential decay. Images are resized to 1000×15001000$\times$ 15001000 × 1500, and rendering uses spherical harmonics of order $l=3$. Progressive densification, opacity pruning, and geometric regularization are applied. Semantic filtering, sampling, and clustering hyperparameters are listed in Table 1.
Table 4: Ablation study on the effect of language-conditioned semantic embedding in our fruit counting pipeline. Each cell reports the predicted count vs. GT, followed by recall percentage. “w/o Language-Conditioned Embedding” disables the CLIP-guided semantic vector $\mathbf{f}_{i}$ during filtering and clustering. “w/ Fruit Field Replacement” replaces our semantic embedding module with the Fruit Field representation and clustering approach from FruitNeRF. Bold values indicate the best result per row.
Dataset
w/ Language-
Conditioned Embedding
w/o Language-
Conditioned Embedding
w/ Fruit Field Replacement
from FruitNeRF
Tree01 173 / 179 (96.6%) 141 / 179 (78.8%) 158 / 179 (88.3%)
Tree02 110 / 113 (97.3%) 93 / 113 (82.3%) 101 / 113 (89.4%)
Tree03 287 / 291 (98.6%) 204 / 291 (70.1%) 251 / 291 (86.3%)
Fuji-SfM 1443 / 1455 (99.2%) 1182 / 1455 (81.2%) 1327 / 1455 (91.2%)
Pfuji-Size_Orch2018 1336 / 1455 (91.8%) 1188 / 1455 (81.6%) 1254 / 1455 (86.2%)
Pfuji-Size_Orch2020 1329 / 1455 (91.3%) 1174 / 1455 (80.7%) 1201 / 1455 (82.5%)
@@ 4.2 Reconstruction Quality and Efficiency
We first evaluate the 3D reconstruction performance of FruitLangGS in terms of both visual fidelity and rendering efficiency. As summarized in Table 2, we compare our method with four baselines: NeRF [16], FruitNeRF [9], Taming3DGS [19], and 3D Gaussian Splatting (3DGS) [17]. Evaluation is performed on nine datasets.
FruitLangGS consistently achieves the highest rendering speed across all datasets. For instance, on Tree01, it reaches 988.86 FPS—over 3× faster than Taming3DGS (277.91 FPS) and nearly 7× faster than 3DGS (144.78 FPS). NeRF and FruitNeRF remain below 0.5 FPS across all cases. On the large-scale Fuji-SfM benchmark, our method maintains 306.88 FPS, significantly outperforming 3DGS (26.63 FPS) and Taming3DGS (196.69 FPS). Notably, even on challenging real-world orchard sequences such as Galafab-West and Schnico-Red-East, our method sustains over 1000 FPS, demonstrating the scalability of our pruning and tile-aware scheduling strategies to complex and cluttered environments.
In terms of reconstruction quality, FruitLangGS achieves results that are competitive with or superior to prior work. On Fuji-SfM, it yields the best SSIM (0.716) and PSNR (21.454), while matching FruitNeRF in LPIPS (0.295). Occlusion-heavy scenes such as Tree03 highlight the advantage of our adaptive Gaussian model: while NeRF and FruitNeRF report SSIM below 0.39 and LPIPS above 0.68, our method maintains SSIM of 0.599 and LPIPS of 0.340. Among the new datasets, Pfuji-Size_Orch2020 demonstrates similar trends—our method achieves the lowest LPIPS (0.227) and highest SSIM (0.749), indicating strong robustness to motion blur and viewpoint variability. These results confirm that FruitLangGS balances reconstruction accuracy with real-time performance, enabling scalable and interactive scene understanding in outdoor orchards.